In 2018, Lisa and Fred, a married couple, have taxable income of $321,000. If they were to file separate tax returns, Lisa would have reported taxable income of $132,000 and Fred would have reported taxable income of $189,000. What is the couple’s marriage benefit?
Answer #1
Tax payable, Filing jointly 66099 =64179+(321000-315000)*32%
Tax payable, Filing single:
Lisa 25969.5 =14089.5+(132000-82500)*24%
Fred 42169.5 =32089.5+(189000-157500)*32%
Total 68139
Couple’s marriage benefit 2040 =68139-66099